# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v3

@@ V Conclusions
Appendix A Proof of Proposition 1
@@
Appendix B ELBO Derivation
@@
$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}{q({\mathbf{x}}_{0})q(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}\right]$ (44c)
@@
where we augment the GP using inducing points and use the variational sparse GP in Eq. (44c); We factorize the joint likelihood into sequential terms in Eq. (44e), where for $t=1$ we have the log-term as $\log p({\mathbf{y}}_{1}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})$. ∎
Appendix C Proof of Corollary 1
Proof.
Similar proof of this corollary has been discussed in the literature; see [42]. For ease of reference, we also detail here. Consider two arbitrary time steps $t$ and $t^{\prime}$, we have
Since ${{\bm{\theta}}}_{t}$ and ${{\bm{\theta}}}_{t^{\prime}}$ are deterministic, the joint distribution of $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is Gaussian with mean $[\bm{\beta}_{t}^{\top},\bm{\beta}_{t^{\prime}}^{\top}]^{\top}$.
The covariance between $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is:
Similarly, the covariance of $\bm{f}({\mathbf{x}}_{t-1})$ with itself is:
@@ Appendix C Proof of Corollary 1
Following the GPSSM literature [37, 31], an alternative derivation of the variational inference algorithm presented in this work can begin with the following ELBO, which explicitly defines the variational distribution over the latent states:
@@
It turns out that $\mathcal{L}_{2}$ is a lower bound of $\mathcal{L}$ in Eq. (22) as summarized in Proposition 1 in the main text. Next, we show how we can use the methodology proposed in [37] to derive the same ELBO as in Eq. (26).
@@
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}).$ (52)
The term $p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0:T}|\mathbf{w},\tilde{f})$ is factorized as:
@@
which is the objective function we have used in our paper. Note that from Eq. (56d) to Eq. (56e), we have made the following two assumptions, which is similar to the work in [37]:
@@
These two approximations used here are similar to those in [37], with the key difference being the focus on the ETGP in this paper rather than the GP. The rationale behind these approximations is as follows:
  • Assumption 1) assumes that the posterior distribution of the latent state at time $t-1$ based on observations up to $t-1$ is nearly identical to that based on observations up to $t$. This holds particularly well when $t$ is large, as the additional observation at time $t$ has minimal impact on the state estimate due to the accumulated information over time.
@@
It can be observed that, by adopting the approach from [37], which relies on two specific assumptions, one can also derive the same objective function as presented in our main text. However, our contribution in this paper lies in providing an alternative perspective—one that is more straightforward and assumption-free—by leveraging only the EnKF for inference. This not only simplifies the derivation but also enhances the practicality of the method.

### v4

@@ V Conclusions
Appendix A Proof of Corollary 1
Proof.
Similar proof of this corollary has been discussed in the literature; see [maronas2022efficient]. For ease of reference, we also detail here. Consider two arbitrary time steps $t$ and $t^{\prime}$, we have
Since the parameters $\bm{\theta}_{t}$ and $\bm{\theta}_{t^{\prime}}$ are deterministic when conditioned on the inputs $\mathbf{x}_{t}$ and $\mathbf{x}_{t^{\prime}}$, the joint distribution of $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is Gaussian with mean $\bm{\beta}_{t}^{\top},\bm{\beta}_{t^{\prime}}^{\top}]^{\top}$. The covariance between $\bm{f}({\mathbf{x}}_{t-1})$ and $\bm{f}({\mathbf{x}}_{t^{\prime}-1})$ is:
Similarly, the covariance of $\bm{f}({\mathbf{x}}_{t-1})$ with itself is:
Appendix B Proof of Proposition 1
@@
Appendix C Proposition 2: ELBO Derivation
@@
$=\mathbb{E}_{q}\left[\log\frac{p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p(\mathbf{w})p(\tilde{f}|{\mathbf{u}})p({\mathbf{u}})}{q({\mathbf{x}}_{0})q(\mathbf{w})p(\tilde{f}|{\mathbf{u}})q({\mathbf{u}})}\right]$ (50c)
@@
[…]
@@
Following the GPSSM literature [lin2023ensemble, ialongo2019overcoming], an alternative derivation of the variational inference algorithm presented in this work can begin with the following ELBO, which explicitly defines the variational distribution over the latent states:
@@
It turns out that $\mathcal{L}_{2}$ is a lower bound of $\mathcal{L}$ in Eq. (22) as summarized in Proposition 1 in the main text. Next, we show how we can use the methodology proposed in [lin2023ensemble] to derive the same ELBO as in Eq. (26).
@@
[…]
@@
which is the objective function we have used in our paper. Note that from Eq. (56d) to Eq. (56e), we have made the following two assumptions, which is similar to the work in [lin2023ensemble]:
@@
These two approximations used here are similar to those in [lin2023ensemble], with the key difference being the focus on the ETGP in this paper rather than the GP. The rationale behind these approximations is as follows:
  • Assumption 1) assumes that the posterior distribution of the latent state at time t1t-1 based on observations up to t1t-1 is nearly identical to that based on observations up to $t$. This holds particularly well when $t$ is large, as the additional observation at time $t$ has minimal impact on the state estimate due to the accumulated information over time.
@@
It can be observed that, by adopting the approach from [lin2023ensemble], which relies on two specific assumptions, one can also derive the same objective function as presented in our main text. However, our contribution in this paper lies in providing an alternative perspective—one that is more straightforward and assumption-free—by leveraging only the EnKF for inference. This not only simplifies the derivation but also enhances the practicality of the method.
Supplement E Implementation Details
Throughout the paper, we use the network architecture and hyperparameters specified in Table III. For further details, please refer to the code repository available online.
TABLE III: Implementation details for experimental comparisons.
Component ETGPSSM AD-EnKF [chen2022autodifferentiable] EnVI [lin2023ensemble]
Transition Model ETGP (1 GP + NN flows) FCN $d_{x}$ independent GPs
FCN Architecture $d_{x}\to 128\to 64\to 2d_{x}$ $d_{x}\to 128\to 64\to d_{x}$
Ensemble Size ($N$) 200 200 200
Optimizer Adam Adam Adam
Learning Rate 0.005 0.005 0.005
Batch Size Full sequence Full sequence Full sequence
Kernel Function SE SE
Activation Function ReLU ReLU
Training Epochs 1000 1000 1000
Early Stopping Yes Yes Yes
In Section IV-B of the main paper, we adopt the following four performance metrics [spantini2022coupling] to collectively provide a comprehensive assessment of both point estimation accuracy and uncertainty quantification in filtering.
  • Root Mean Squared Error (RMSE): Captures the average deviation between the estimated states and the true reference states.
  • Spread: Defined as the root mean trace of the ensemble covariance, this metric reflects the internal variability of the ensemble.
  • Coverage Probability: Measures how frequently the true states fall within the 95% empirical confidence intervals (constructed from the 2.5% and 97.5% quantiles) of the ensemble marginals.
  • Continuous Ranked Probability Score (CRPS): A comprehensive scoring rule that quantifies how well the predictive distribution matches the observed outcomes [gneiting2007probabilistic]. It evaluates the discrepancy between the forecasted and actual cumulative distribution functions, with lower values signifying more accurate predictions.